OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -3 f(1) = 6 f(n) = -2*f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function?

OpenStudy (anonymous):

(A) -3, 6, -9, -12, … (B) -3, 20, -95, 480, … (C) -3, 6, -9, 12, … (D) -3, -20, -95, -480, …

OpenStudy (anonymous):

I know that the answers cannot be B and D

OpenStudy (anonymous):

you can toss out B and D

OpenStudy (anonymous):

Yup that's what I said

OpenStudy (anonymous):

\[f(2)=-2f(1)-f(0)\\ f(2)=-2\times 6-(-3)\\ f(2)=-12+3=-9\]

OpenStudy (anonymous):

but both A and C have \(-9\) as the third number, so you need to compute \(f(3)\)

OpenStudy (anonymous):

i.e. you need to compute \[f(3)=-2\times f(2)-f(1)\]

OpenStudy (anonymous):

f(3) = -2 * -9 - 6

OpenStudy (anonymous):

f(3) = 12

OpenStudy (anonymous):

So the answer is C

OpenStudy (anonymous):

Right, answer is C?

OpenStudy (anonymous):

yeah it is always C

OpenStudy (anonymous):

Thank you so much!!